Dinitrogen oxide can be produced by heating solid ammonium nitrate to 200 degrees C or more. (At 200.0 degrees C, N2O and H2O are both gases.) If the 2.00 kg of solid NH4NO3 were heated to 200.0 degrees C in a 250. L steel reaction vessel, what would the pressure in the vessel be following the reaction? Assume the reaction goes to completion.
1
Expert's answer
2015-11-30T12:05:11-0500
Solution:
NH4NO3 → N2O↑ + 2H2O↑ (t = 200˚C)
P - ?
PV = nRT,
so P = nRT / V
R = 0.0821 L*atm/mol*K (gas constant)
V = 250 L
T = t˚ + 273K = 200˚C + 273 K = 473 K
n = m / M
M (NH4NO3) = 80 g/mol
n (NH4NO3) = 2000 g / 80 g/mol = 25 mol
P = (25mol * 0.0821 L*atm/mol*K * 473K) / 250 L = 3.89 atm